OpenStudy (anonymous):

Consider the function f(x) = x 3 and use the Newton method to analytically find expressions for the sequence members xn when you start from the initial condition x0 = 2. Show that the Newton method converges towards the correct solution. How often do you have to iterate to get within 0.01 of the correct solution? @ganeshie8

OpenStudy (anonymous):

So far I have gotten that f'(x) = 3x^2 so Xn+1 = Xn - Xn^3/ 3*Xn^2 which is the same as Xn+1 = Xn - Xn/3 So when X0 = 2 X1 = 2 - 2/3 = 4/3 X2 = 4/3 - (4/3)/3 = 8/9 X3 = 8/9 - (8/9) /3 = 16/27 My Question is how do I show that it converges towards the correct solution and how often do you have to iterate to get within 0.01 of the correct solution?